Marie's office cubicle measures 10 ft by 12 ft. Write an equation to find the amount (x) that should be added to both the current length and width to make the area twice the size it is now. Solve the equation. What will the new dimension be?

Answer by ankor@dixie-net.com(22740)

Marie's office cubicle measures 10 ft by 12 ft.
Write an equation to find the amount (x) that should be added to both the current length and width to make the area twice the size it is now
Solve the equation. What will the new dimension be?
:
Find the area now: 10 * 12 = 120 sq/ft, double the area = 240 sq/ft
:
(x+10)(x+12) = 240
FOIL
x^2 + 12x + 10x + 120 = 240
x^2 + 22x + 120 - 240 = 0
x^2 + 22x - 120 = 0; is the equation
Won't factor, use the quadratic formula to find x, I got a positive solution of
x ~ 4.524 ft or 4 ft 6.3 inches to add to each dimension